OpenStudy (anonymous):

Word Problem: Sarah has 20 quarters and nickels. She has a total of three dollars. How many nickels does she have?

OpenStudy (anonymous):

.25x+.1y=3 x+y=20

OpenStudy (anonymous):

x denote quater while y denotes dimee.

OpenStudy (anonymous):

oops, not dime .25x+.05y=3 x+y=20 x denote quater while y denotes nickel.

OpenStudy (anonymous):

solve for x and y

OpenStudy (anonymous):

HA! its ten, ten nickels and ten quarters make 3 dollars with a total of 20 coins... BAM

OpenStudy (anonymous):

thanks so much.

OpenStudy (anonymous):

i assume a quarter is 25 cents

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

gah, i couldnt see any of iranmeah91's answers when i was posting.

OpenStudy (anonymous):

What do they call .25 cent in Australia?

OpenStudy (anonymous):

the closest we have is a 20 cent piece goes 5c coin 10c coin 20c coin 50c coin 1dollar coin 2dollar coin 5dollar note 10dollar note 20dollar note 50 dollar note 100 dollar note not sure about any higher values if we have them, they aren't in common use and we dont have any special names for any of them except nicknames (like 10 dollar note is a bluey, or 20 is a lobster, 50 is a pineapple all because of the colour)

OpenStudy (anonymous):

you're actually funny lol

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

would you be willing to do another problem?

OpenStudy (anonymous):

i can try

OpenStudy (anonymous):

solve for "x" : fh/2x= mn

OpenStudy (anonymous):

does that mean you have 5 unknowns?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

what is fh and mn?

OpenStudy (anonymous):

just variables, we are only solving for "x".

OpenStudy (anonymous):

you can still have variables in your answer.

OpenStudy (anonymous):

yeah, but if you need to solve for x and you dont know if fh and mn are constants or variable, you need to change how you ansewr the question

OpenStudy (anonymous):

ok fh/2x=mn fh/mn=2x (fh/mn)/2=x

OpenStudy (anonymous):

hmm let me see.. i think that's it

OpenStudy (anonymous):

yea thanks

OpenStudy (anonymous):

yay us, we can do algebra!!!

OpenStudy (anonymous):

lol yep, we did it